I'm having difficult time doing word problems for applications of linear equations.
The number of pages in the U.S. tax code exceeds the number of pages in the King James Bible by 18,528. The combined number of pages in the U.S. tax code and the King James Bible is 21,472. Determine the number of pages in each.
Thanks for any help you can give me.

Answer by stanbon(75887)

The number of pages in the U.S. tax code exceeds the number of pages in the King James Bible by 18,528. The combined number of pages in the U.S. tax code and the King James Bible is 21,472. Determine the number of pages in each.
-----------------------------------
t = b + 18528
t + b = 21472
-----------------
Rearrange:
t - b = 18528
t + b = 21472
------------------
Add the two equations to solve for "t":
2t = 40000
t = 20,000 (# of pages in the tax code)
---
Substitute to solve for "b":
20,000 + b = 21472
b = 1472 (# of pages in the bible)
==========================================
Cheers,
Stan H.